OpenStudy (anonymous):

You draw 1 card at random from a well-shuffled deck of 52 cards, set it aside without replacing it, and draw another card from the deck at random. The events "get the king of spades on the first draw" and "get the queen of diamonds on the second draw" are independent events. True or False???

OpenStudy (anonymous):

In probability theory, independence means that the occurrence of one does not affect the probability of the other. In this case there are two possible alternatives, -you could draw queen of diamonds on the first card (or not), which would affect the outcome of the second choice. In which case the first event would directly impact the probability of the second event occuring, so I would guess that these event are not independent. That's how I would take it, maybe someone else could clarify?

OpenStudy (anonymous):

so r u saying that its false

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

I would say false.

OpenStudy (ybarrap):

f P(A|B) = P(A), then the events are independent Let A = getting KS and B = getting QD You would not expect P(B) = P(B on 2nd draw|A on 1st draw), since P(B) = 1/52, but P(B|A) = 1/51

OpenStudy (anonymous):

so... u say true?

OpenStudy (ybarrap):

False, because the second event is dependent of the 1st event.

OpenStudy (anonymous):

oh, right.

OpenStudy (ybarrap):

If you draw a QD on 1st then P(B on second) = zero, otherwise, it's 1/51

OpenStudy (anonymous):

ok.

OpenStudy (ybarrap):

So P(B on second) is totally dependent on what happens on the 1st draw

OpenStudy (ybarrap):

True now. Because now P(B on second) is totally independent of what happened on the 1st draw.

OpenStudy (anonymous):

True

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

how do u do this? :/

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

The intersection \(\cap\) means an 'and' operator. So you're being asked which hand is in F and in S F is "all the cards in your hand are the same suit" and S is "all the numbers in your card form an uninterrupted sequence" Do you play poker? :P

OpenStudy (anonymous):

no :p

OpenStudy (anonymous):

so how can I solve this

terenzreignz (terenzreignz):

Well, that's a pity. You'll find that the set F is the set of all hands which are 'flushes' while the set S is the set of all hands which are "straight's" In effect, the set \(\large F \cap S\) is the set of all straight flushes...

terenzreignz (terenzreignz):

So, which of those hands have cards which all have the same suit?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

A

OpenStudy (anonymous):

D

terenzreignz (terenzreignz):

It had to be in that order? Besides, C are all spades...

OpenStudy (anonymous):

no it doesn't have to be in that order... C is???

terenzreignz (terenzreignz):

C also consists of a hand that are all the same suit... they're all spades...

OpenStudy (anonymous):

what abt the Jack?

terenzreignz (terenzreignz):

If you actually read your question, the J = 11, Q = 12, K = 13 :P

OpenStudy (anonymous):

yeah I know I mean to say that its a ♥ though

OpenStudy (anonymous):

@ybarrap ?

terenzreignz (terenzreignz):

Crud... my eyes fail me....

terenzreignz (terenzreignz):

Apologies, @music101

terenzreignz (terenzreignz):

They really should colour these things :P

OpenStudy (anonymous):

oh its fine

OpenStudy (ybarrap):

A. satisfies both events, suit and uninterrupted sequence. B. does not because there is a suit ther mixed with the hearts that doesn't look like any suit. Why the rest don't satisfy the conditions is obvious.

terenzreignz (terenzreignz):

So, B, A, and D all have five cards with the same suit... now, among these three, which form an uninterrupted sequence?

OpenStudy (anonymous):

what do u mean an uninterrupted sequence

terenzreignz (terenzreignz):

A straight... err for example, A,2,3,4,5 or 8,9,10,J,Q You get the idea :)

OpenStudy (anonymous):

so A&B?

terenzreignz (terenzreignz):

That is correct :)

OpenStudy (anonymous):

ok :) hope that's correct.. oh btw what did @ ybarrap mean "B is not"